# Perfect state transfer between real pure states

### v1

@@ 5 Perfect state transfer
If $|\sigma_{\mathbf{x}}(M)|=2$, then the same proof used in Theorem 26 works. Now, let $|\sigma_{\mathbf{x}}(M)|\geq 3$ and suppose $\mathbf{x}=\sum_{j\in\sigma_{\mathbf{x}}(M)}\mathbf{u}_{j}$, where $\mathbf{u}_{j}$ is a real eigenvector associated with the eigenvalue $\lambda_{j}\in\sigma_{\mathbf{x}}(M)$. Since $\mathbf{x}$ is periodic, Theorem 8 implies that we may write $\frac{\lambda_{1}-\lambda_{j}}{\lambda_{1}-\lambda_{2}}=\frac{p_{j}}{q_{j}}$ for each $j\geq 3$, where $p_{j},q_{j}$ are integers such that $\operatorname{gcd}(p_{j},q_{j})=1$. If each $p_{j}$ and $q_{j}$ are odd, then define $\mathbf{y}:=\mathbf{u}_{1}-\sum_{j\neq 1}\mathbf{u}_{j}$. Since $|\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)|=1$, invoking Theorem 23(2a) yields PST between $\mathbf{x}$ and $\mathbf{y}$. If some $p_{j}$s are even and each $q_{j}$ is odd, then let $\sigma_{1}=\{\lambda_{j}:p_{j}\ \text{is even}\}$, $\sigma_{2}=\sigma_{\mathbf{x}}(M)\backslash\sigma_{1}$, and define $\mathbf{y}:=\sum_{j\in\sigma_{1}}\mathbf{u}_{j}-\sum_{j\in\sigma_{2}}\mathbf{u}_{j}$. Note that $\lambda_{1}\in\sigma_{1}$, $\lambda_{2}\in\sigma_{2}$ and $\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)=\sigma_{1}$ has at least two elements. Since $\lambda_{1}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$ and $\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$, applying Theorem 23(2c) yields PST between $\mathbf{x}$ and $\mathbf{y}$. Finally, if some $q_{j}$s are even, then let $\sigma_{2}:=\{\lambda_{\ell}:\nu_{2}(q_{\ell}))=\eta\}$, where $\eta=\max_{j\geq 3}\nu_{2}(q_{j})>0$. Define $\mathbf{y}=\sum_{j\in\sigma_{1}}\mathbf{u}_{j}-\sum_{j\in\sigma_{2}}\mathbf{u}_{j}$, where $\sigma_{1}=\sigma_{\mathbf{x}}(M)\backslash\sigma_{2}$. Since $\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)=\sigma_{1}$ has at least two elements and $\lambda_{1},\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$, Theorem 23(2b) again yields PST between $\mathbf{x}$ and $\mathbf{y}$. In all cases, $\mathbf{y}$ is unique by Lemma 22(3). Moreover, $\mathbf{y}$ is real because the $\mathbf{u}_{j}$s are all real. ∎
@@
Let $\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}$ with $\mathbf{y}\neq\pm\mathbf{x}$. For all $\tau>0$ and for all integers $m_{1},m_{2}\geq 1$ such that $m_{1}+m_{2}\leq n$, there exists a real symmetric matrix $M$ such that perfect state transfer occurs between $\mathbf{x}$ and $\mathbf{y}$ relative to $M$ at time $\tau$, $|\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)|=m_{1}$ and $|\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)|=m_{2}$.
@@
$M=\sum_{j=1}^{n}\theta_{j}\mathbf{w}_{j}\mathbf{w}_{j}^{T}$
@@
where each $g_{j}=\operatorname{gcd}(b_{j},b_{2})$. Thus, $q=\operatorname{lcm}(q_{3},\ldots,q_{m_{1}+m_{2}})=\operatorname{lcm}(\frac{b_{2}}{g_{3}},\ldots,\frac{b_{2}}{g_{m_{1}+m_{2}}})=\frac{b_{2}}{g}$. We have two cases. First, if $m_{1}=1$, then each $p_{j}$ and $q_{j}$ is odd. In this case, invoking Theorem 23(2a) yields PST between $\mathbf{x}$ and $\mathbf{y}$ relative to $M$ at time $\frac{\pi q}{\theta_{1}-\theta_{2}}=\frac{\pi b_{2}/g}{\pi b_{2}/(g\tau)}=\tau$. Now, if $m_{1}\geq 2$, then $b_{\ell}$ is odd for each $\theta_{\ell}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$, and so $\nu_{2}(g_{\ell})=\nu_{2}(\operatorname{gcd}(b_{\ell},b_{2}))=0$. Consequently, $\nu_{2}(q_{\ell})=\nu_{2}(\frac{b_{2}}{g_{\ell}})=\nu_{2}(b_{2})$ for each $\theta_{\ell}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$. Moreover, $b_{h}$ is even for each $\theta_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$, and so $\nu_{2}(g_{h})=\nu_{2}(\operatorname{gcd}(b_{h},b_{2}))\geq 1$. Consequently, $\nu_{2}(q_{h})=\nu_{2}(\frac{b_{2}}{g_{h}})<\nu_{2}(b_{2})$ for each $\theta_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$. Equivalently,
@@
for all $\lambda_{\ell},\lambda_{k}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$ and $\lambda_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)\backslash\{\lambda_{1},\lambda_{2}\}$, where each $\nu_{2}(q_{h})$ above is absent whenever $|\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)|=2$. Finally, invoking Theorem 23(2b) yields PST between $\mathbf{x}$ and $\mathbf{y}$ at time $\tau$.
@@ 6 Complete graphs and cycles
For $K_{n}$, $U_{A}(t)=e^{-it}\left((e^{itn}-1)\frac{1}{n}J+I\right).$ Thus, $e^{it}U_{A}(t)\mathbf{x}=\mathbf{x}+\left(\frac{(e^{itn}-1){\bf{1}}^{\top}\mathbf{x}}{n}\right){\bf{1}}.$ If $\mathbf{x}=\frac{1}{n}{\bf{1}}$ or $\mathbf{x}\perp{\bf{1}}$, then $\mathbf{x}$ is fixed. Otherwise, $e^{i\pi/n}U_{A}(\tau)\mathbf{x}=\mathbf{x}-\frac{2({\bf{1}}^{\top}\mathbf{x})}{n}{\bf{1}}$, with $\frac{\pi}{n}$ as the minimum PST time by Theorem 28. ∎
@@
For cycles, we adopt the convention that the vertices of $C_{n}$ are labelled so that vertices $j,k$ are adjacent whenever $|k-j|\equiv 1\mod n.$ The eigenvalues and eigenvectors of $C_{n}$ are well-known, see [brouwer2011spectra, Section 1.4.3]. For our purposes, we provide normalized eigenvectors for $C_{n}$ in the following lemma.
@@
    $n=12m$, $\mathbf{x}=a(\alpha_{1}\mathbf{v}_{3m}+\alpha_{2}\mathbf{v}_{9m})+b(\beta_{1}\mathbf{v}_{m}+b\beta_{2}\mathbf{v}_{11m})+c(\gamma_{1}\mathbf{v}_{5m}+\gamma_{2}\mathbf{v}_{7m})$ and $\mathbf{y}=a(\alpha_{1}\mathbf{v}_{m}+\alpha_{2}\mathbf{v}_{9m})-b(\beta_{1}\mathbf{v}_{m}+\beta_{2}\mathbf{v}_{11m})-c(\gamma_{1}\mathbf{v}_{5m}+\gamma_{2}\mathbf{v}_{7m})$. In this case, $\sigma_{\mathbf{x}}(M)=\{0,\pm\sqrt{3}\}$.
@@
    $n=8m$, $\mathbf{x}=a(\alpha_{1}\mathbf{v}_{2m}+\alpha_{2}\mathbf{v}_{6m})+b(\beta_{1}\mathbf{v}_{m}+\beta_{2}\mathbf{v}_{7m})+c(\gamma_{1}\mathbf{v}_{3m}+\gamma_{2}\mathbf{v}_{5m})$ and $\mathbf{x}=a(\alpha_{1}\mathbf{v}_{2m}+\alpha_{2}\mathbf{v}_{6m})-b(\beta_{1}\mathbf{v}_{m}+\beta_{2}\mathbf{v}_{7m})-c(\gamma_{1}\mathbf{v}_{3m}+\gamma_{2}\mathbf{v}_{5m})$. In this case $\sigma_{\mathbf{x}}(M)=\{0,\pm\sqrt{2}\}$.
In all cases above, $a,b,c,d,e\in\mathbb{R}$ are such that $a^{2}+b^{2}+c^{2}=\|x\|^{2}$ in (2)-(4) with $a,b,c\neq 0$, and $a^{2}+b^{2}+c^{2}+d^{2}+e^{2}=\|x\|^{2}$ otherwise. Moreover, $(\alpha_{1},\alpha_{2}),(\beta_{1},\beta_{2}),(\gamma_{1},\gamma_{2})\in\mathbb{R}^{2}\backslash\{(0,0)\}$ such that $\alpha_{1}^{2}+\alpha_{2}^{2}=\beta_{1}^{2}+\beta_{2}^{2}=\gamma_{1}^{2}+\gamma_{2}^{2}=1$. The minimum PST time in (1)-(4) is $\pi$, $\frac{\pi}{2}$, $\frac{\pi}{\sqrt{3}}$ and $\frac{\pi}{\sqrt{2}}$, respectively.
@@
Case 2. Let $\sigma_{\mathbf{x}}(A)\subseteq\{\frac{a}{2},\frac{1}{2}(a\pm b_{1}\sqrt{\Delta}),\frac{1}{2}(a\pm b_{2}\sqrt{\Delta})\}$, where $a,b_{1},b_{2},\Delta$ are integers such that $b_{1}>b_{2}>0$ and $\Delta>1$ is square-free. Note that $\frac{a}{2}$ is an integer as it is a root of a polynomial with integer coefficients. Since $|\frac{a}{2}|\leq 2$, we get $a\in\{0,\pm 2,\pm 4\}$. If $a=2$, then $|\frac{1}{2}(2\pm b_{j}\sqrt{\Delta})|\leq 2$ implies that $b_{j}=1$ and $\Delta\in\{2,3\}$. However, $\frac{1}{2}(2\pm\sqrt{\Delta})$ are not quadratic integers whenever $\Delta\in\{2,3\}$. Thus, $a\neq 2$ and similarly, $a\neq-2$. Now, if $a=0$, then $|\frac{1}{2}b_{j}\sqrt{\Delta}|\leq 2$, where $\frac{1}{2}b_{j}$ is an integer, and so we get that $\frac{1}{2}b_{j}=1$ and $\Delta\in\{2,3\}$. Since $b_{1}>b_{2}$ and $\sigma_{\mathbf{x}}(A)$ is closed under algebraic conjugates, $|\sigma_{\mathbf{x}}(A)|\in\{4,5\}$ cannot hapen. Thus, $|\sigma_{\mathbf{x}}(A)|=3$. We have the following subcases.
  • Let $\Delta=3$. Then $\sigma_{\mathbf{x}}(A)=\{0,\pm\sqrt{3}\}$. By Lemma 34, $n=12m$ and $j\in\{\frac{n}{4},\frac{n}{12},\frac{5n}{12}\}$. Thus, $\mathbf{x}=a\mathbf{z}+b\mathbf{u}+c\mathbf{w}$, where $\mathbf{z}=\alpha_{1}\mathbf{v}_{\frac{n}{4}}+\alpha_{2}\mathbf{v}_{\frac{3n}{4}}$, $\mathbf{u}=\beta_{1}\mathbf{v}_{\frac{n}{12}}+\beta_{2}\mathbf{v}_{\frac{11n}{12}}$ and $\mathbf{w}=\gamma_{1}\mathbf{v}_{\frac{5n}{12}}+\gamma_{2}\mathbf{v}_{\frac{7n}{12}}$.
@@
Let $\mathbf{x}=\sum_{j=0}^{m-1}\mathbf{e}_{4j+1}$ and $\mathbf{y}=\sum_{j=0}^{m-1}\mathbf{e}_{4j+3}$ in $C_{4m}$. We may write $\mathbf{x}=\frac{\sqrt{n}}{2}(\frac{1}{\sqrt{2}}\mathbf{v}_{m}+\mathbf{v}_{0}+\mathbf{v}_{2m})$ and $\mathbf{y}=-\frac{\sqrt{n}}{2}(\frac{1}{\sqrt{2}}\mathbf{v}_{m}-\mathbf{v}_{0}-\mathbf{v}_{2m})$. Invoking Theorem 35(2) with $a=c$, $\alpha_{1}=1$ and $b=\frac{1}{\sqrt{2}}$ yields PST between $\mathbf{x}$ and $\mathbf{y}$ at time $\frac{\pi}{2}$. In particular, if $m=2$, then we recover the fact that $C_{8}$ admits plus PST between $\mathbf{x}=\mathbf{e}_{1}+\mathbf{e}_{5}$ and $\mathbf{y}=\mathbf{e}_{3}+\mathbf{e}_{7}$ [kim2024generalization]. If $m=3$, then $C_{12}$ admits PST between $\mathbf{x}=\mathbf{e}_{1}+\mathbf{e}_{5}+\mathbf{e}_{9}$ and $\mathbf{y}=\mathbf{e}_{3}+\mathbf{e}_{7}+\mathbf{e}_{11}$. This complements the fact that $C_{12}$ does not admit $s$-pair PST.

### v2

@@ 5 Perfect state transfer
If $|\sigma_{\mathbf{x}}(M)|=2$, then the same proof used in Theorem 26 works. Now, let $|\sigma_{\mathbf{x}}(M)|\geq 3$ and suppose $\mathbf{x}=\sum_{j\in\sigma_{\mathbf{x}}(M)}\mathbf{u}_{j}$, where $\mathbf{u}_{j}$ is a real eigenvector associated with the eigenvalue $\lambda_{j}\in\sigma_{\mathbf{x}}(M)$. Since $\mathbf{x}$ is periodic, Theorem 8 implies that we may write $\frac{\lambda_{1}-\lambda_{j}}{\lambda_{1}-\lambda_{2}}=\frac{p_{j}}{q_{j}}$ for each $j\geq 3$, where $p_{j},q_{j}$ are integers such that $\operatorname{gcd}(p_{j},q_{j})=1$. If each $p_{j}$ and $q_{j}$ are odd, then define $\mathbf{y}:=\mathbf{u}_{1}-\sum_{j\neq 1}\mathbf{u}_{j}$. Since $|\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)|=1$, invoking Theorem 23(2b) yields PST between $\mathbf{x}$ and $\mathbf{y}$. If some $p_{j}$’s are even and each $q_{j}$ is odd, then let $\sigma_{1}=\{\lambda_{j}:p_{j}\ \text{is even}\}$, $\sigma_{2}=\sigma_{\mathbf{x}}(M)\backslash\sigma_{1}$, and define $\mathbf{y}:=\sum_{j\in\sigma_{1}}\mathbf{u}_{j}-\sum_{j\in\sigma_{2}}\mathbf{u}_{j}$. Note that $\lambda_{1}\in\sigma_{1}$, $\lambda_{2}\in\sigma_{2}$ and $\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)=\sigma_{1}$ has at least two elements. Since $\lambda_{1}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$ and $\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$, applying Theorem 23(2b) yields PST between $\mathbf{x}$ and $\mathbf{y}$. Finally, if some $q_{j}$’s are even, then let $\sigma_{2}:=\{\lambda_{\ell}:\nu_{2}(q_{\ell}))=\eta\}$, where $\eta=\max_{j\geq 3}\nu_{2}(q_{j})>0$. Define $\mathbf{y}=\sum_{j\in\sigma_{1}}\mathbf{u}_{j}-\sum_{j\in\sigma_{2}}\mathbf{u}_{j}$, where $\sigma_{1}=\sigma_{\mathbf{x}}(M)\backslash\sigma_{2}$. Since $\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)=\sigma_{1}$ has at least two elements and $\lambda_{1},\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$, Theorem 23(2a) again yields PST between $\mathbf{x}$ and $\mathbf{y}$. In all cases, $\mathbf{y}$ is unique by Lemma 22(3). Moreover, $\mathbf{y}$ is real because the $\mathbf{u}_{j}$s are all real. ∎
@@
Let $\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}\backslash\{{\bf{0}}\}$ with $\mathbf{y}\neq\pm\mathbf{x}$. For all $\tau>0$ and for all integers $m_{1},m_{2}\geq 1$ such that $m_{1}+m_{2}\leq n$, there exists a real symmetric matrix $M$ such that perfect state transfer occurs between $\mathbf{x}$ and $\mathbf{y}$ relative to $M$ at time $\tau$, $|\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)|=m_{1}$ and $|\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)|=m_{2}$.
@@
$M=\sum_{j=1}^{n}\theta_{j}\mathbf{w}_{j}\mathbf{w}_{j}^{\top}$
@@
where each $g_{j}=\operatorname{gcd}(b_{j},b_{2})$. Thus, $q=\operatorname{lcm}(q_{3},\ldots,q_{m_{1}+m_{2}})=\operatorname{lcm}(\frac{b_{2}}{g_{3}},\ldots,\frac{b_{2}}{g_{m_{1}+m_{2}}})=\frac{b_{2}}{g}$. We have two cases. First, if $m_{1}=1$, then each $p_{j}$ and $q_{j}$ is odd. In this case, invoking Theorem 23(2b) yields PST between $\mathbf{x}$ and $\mathbf{y}$ relative to $M$ at time $\frac{\pi q}{\theta_{1}-\theta_{2}}=\frac{\pi b_{2}/g}{\pi b_{2}/(g\tau)}=\tau$. Now, if $m_{1}\geq 2$, then $b_{\ell}$ is odd for each $\theta_{\ell}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$, and so $\nu_{2}(g_{\ell})=\nu_{2}(\operatorname{gcd}(b_{\ell},b_{2}))=0$. Consequently, $\nu_{2}(q_{\ell})=\nu_{2}(\frac{b_{2}}{g_{\ell}})=\nu_{2}(b_{2})$ for each $\theta_{\ell}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$. Moreover, $b_{h}$ is even for each $\theta_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$, and so $\nu_{2}(g_{h})=\nu_{2}(\operatorname{gcd}(b_{h},b_{2}))\geq 1$. Consequently, $\nu_{2}(q_{h})=\nu_{2}(\frac{b_{2}}{g_{h}})<\nu_{2}(b_{2})$ for each $\theta_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$. Equivalently,
@@
for all $\theta_{\ell},\theta_{k}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$ and $\theta_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)\backslash\{\theta_{1},\theta_{2}\}$, where each $\nu_{2}(q_{h})$ above is absent whenever $|\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)|=2$. Finally, invoking Theorem 23(2a) yields PST between $\mathbf{x}$ and $\mathbf{y}$ at time $\tau$.
@@ 6 Complete graphs and cycles
For $K_{n}$, $U_{A}(t)=e^{-it}\left((e^{itn}-1)\frac{1}{n}J+I\right).$ Thus, $e^{it}U_{A}(t)\mathbf{x}=\mathbf{x}+\left(\frac{(e^{itn}-1){\bf{1}}^{\top}\mathbf{x}}{n}\right){\bf{1}}.$ If $\mathbf{x}=\frac{1}{n}{\bf{1}}$ or $\mathbf{x}\perp{\bf{1}}$, then $\mathbf{x}$ is fixed. Otherwise, $\mathbf{x}$ is periodic with $\rho=\frac{2\pi}{n}$. Invoking Corollary 28, $\frac{\pi}{n}$ is the minimum PST time from $\mathbf{x}$, in which case we obtain $e^{i\pi/n}U_{A}(\tau)\mathbf{x}=\mathbf{x}-\frac{2({\bf{1}}^{\top}\mathbf{x})}{n}{\bf{1}}$. ∎
@@
For cycles $C_{n}$, we adopt the convention that $V(C_{n})=\mathbb{Z}_{n}$ where vertices $j,k$ are adjacent if and only if $|k-j|\equiv 1\mod n.$ The eigenvalues and eigenvectors of $C_{n}$ are well-known, see [brouwer2011spectra, Section 1.4.3]. For our purposes, we provide normalized eigenvectors for $C_{n}$ in the following lemma.
@@
    $n=12m$, $\mathbf{x}=a(\alpha_{1}\mathbf{v}_{3m}+\alpha_{2}\mathbf{v}_{9m})+b(\beta_{1}\mathbf{v}_{m}+\beta_{2}\mathbf{v}_{11m})+c(\gamma_{1}\mathbf{v}_{5m}+\gamma_{2}\mathbf{v}_{7m})$ and $\mathbf{y}=a(\alpha_{1}\mathbf{v}_{3m}+\alpha_{2}\mathbf{v}_{9m})-b(\beta_{1}\mathbf{v}_{m}+\beta_{2}\mathbf{v}_{11m})-c(\gamma_{1}\mathbf{v}_{5m}+\gamma_{2}\mathbf{v}_{7m})$. In this case, $\sigma_{\mathbf{x}}(M)=\{0,\pm\sqrt{3}\}$.
@@
    $n=8m$, $\mathbf{x}=a(\alpha_{1}\mathbf{v}_{2m}+\alpha_{2}\mathbf{v}_{6m})+b(\beta_{1}\mathbf{v}_{m}+\beta_{2}\mathbf{v}_{7m})+c(\gamma_{1}\mathbf{v}_{3m}+\gamma_{2}\mathbf{v}_{5m})$ and $\mathbf{y}=a(\alpha_{1}\mathbf{v}_{2m}+\alpha_{2}\mathbf{v}_{6m})-b(\beta_{1}\mathbf{v}_{m}+\beta_{2}\mathbf{v}_{7m})-c(\gamma_{1}\mathbf{v}_{3m}+\gamma_{2}\mathbf{v}_{5m})$. In this case $\sigma_{\mathbf{x}}(M)=\{0,\pm\sqrt{2}\}$.
In all cases above, $a,b,c,d,e\in\mathbb{R}$ are such that $a^{2}+b^{2}+c^{2}=\|\mathbf{x}\|^{2}$ in (2)-(4) with $a,b,c\neq 0$, and $a^{2}+b^{2}+c^{2}+d^{2}+e^{2}=\|\mathbf{x}\|^{2}$ otherwise. Moreover, $(\alpha_{1},\alpha_{2}),(\beta_{1},\beta_{2}),(\gamma_{1},\gamma_{2})\in\mathbb{R}^{2}\backslash\{(0,0)\}$ such that $\alpha_{1}^{2}+\alpha_{2}^{2}=\beta_{1}^{2}+\beta_{2}^{2}=\gamma_{1}^{2}+\gamma_{2}^{2}=1$. The minimum PST time in (1)-(4) is $\pi$, $\frac{\pi}{2}$, $\frac{\pi}{\sqrt{3}}$ and $\frac{\pi}{\sqrt{2}}$, respectively.
@@
Case 2. Let $\sigma_{\mathbf{x}}(A)\subseteq\{\frac{a}{2},\frac{1}{2}(a\pm b_{1}\sqrt{\Delta}),\frac{1}{2}(a\pm b_{2}\sqrt{\Delta})\}$, where $a,b_{1},b_{2},\Delta$ are integers such that $b_{1}>b_{2}>0$ and $\Delta>1$ is square-free. Note that $\frac{a}{2}$ is an integer as it is a root of a polynomial with integer coefficients. Since $|\frac{a}{2}|\leq 2$, we get $a\in\{0,\pm 2,\pm 4\}$. If $a=2$, then $|\frac{1}{2}(2\pm b_{j}\sqrt{\Delta})|\leq 2$ implies that $b_{j}=1$ and $\Delta\in\{2,3\}$. However, $\frac{1}{2}(2\pm\sqrt{\Delta})$ are not quadratic integers whenever $\Delta\in\{2,3\}$. Thus, $a\neq 2$ and similarly, $a\neq-2$. Now, if $a=0$, then $|\frac{1}{2}b_{j}\sqrt{\Delta}|\leq 2$, where $\frac{1}{2}b_{j}$ is an integer, so we get $\frac{1}{2}b_{j}=1$ and $\Delta\in\{2,3\}$. Since $b_{1}>b_{2}$ and $\sigma_{\mathbf{x}}(A)$ is closed under algebraic conjugates, $|\sigma_{\mathbf{x}}(A)|\in\{4,5\}$ cannot hapen. Thus, $|\sigma_{\mathbf{x}}(A)|=3$. We have the following subcases.
  • Let $\Delta=3$. Then $\sigma_{\mathbf{x}}(A)=\{0,\pm\sqrt{3}\}$. By Lemma 34, $n=12m$ and $j\in\{\frac{n}{4},\frac{n}{12},\frac{5n}{12}\}$. Therefore, $\mathbf{x}=a\mathbf{z}+b\mathbf{u}+c\mathbf{w}$, where $\mathbf{z}=\alpha_{1}\mathbf{v}_{\frac{n}{4}}+\alpha_{2}\mathbf{v}_{\frac{3n}{4}}$, $\mathbf{u}=\beta_{1}\mathbf{v}_{\frac{n}{12}}+\beta_{2}\mathbf{v}_{\frac{11n}{12}}$ and $\mathbf{w}=\gamma_{1}\mathbf{v}_{\frac{5n}{12}}+\gamma_{2}\mathbf{v}_{\frac{7n}{12}}$.
@@
Let $\mathbf{x}=\sum_{j=0}^{m-1}\mathbf{e}_{4j}$ and $\mathbf{y}=\sum_{j=0}^{m-1}\mathbf{e}_{4j+2}$ in $C_{4m}$. We may write $\mathbf{x}=\frac{\sqrt{n}}{4}(\mathbf{v}_{0}+\sqrt{2}\mathbf{v}_{m}+\mathbf{v}_{2m})$ and $\frac{\sqrt{n}}{4}(\mathbf{v}_{0}-\sqrt{2}\mathbf{v}_{m}+\mathbf{v}_{2m})$. Invoking Theorem 35(2) with $a=c=\frac{\sqrt{n}}{4}$, $\beta_{1}=1$ (so that $\beta_{2}=0$) and $b=\frac{\sqrt{n}}{\sqrt{2}}$ yields PST between $\mathbf{x}$ and $\mathbf{y}$ at time $\frac{\pi}{2}$. In particular, if $m=2$, then we recover the fact that $C_{8}$ admits plus PST between $\mathbf{x}=\mathbf{e}_{0}+\mathbf{e}_{4}$ and $\mathbf{y}=\mathbf{e}_{2}+\mathbf{e}_{6}$ [kim2024generalization]. If $m=3$, then $C_{12}$ admits PST between $\mathbf{x}=\mathbf{e}_{0}+\mathbf{e}_{4}+\mathbf{e}_{8}$ and $\mathbf{y}=\mathbf{e}_{2}+\mathbf{e}_{6}+\mathbf{e}_{10}$. This complements the fact that $C_{12}$ does not admit $s$-pair PST.